OpenStudy (anonymous):

the legs of a right triangle are 3 and 4. find the length of the hypotenuse.

OpenStudy (anonymous):

write down the general formula which is :\[c^2 = x^2 + y^2\]

OpenStudy (xkehaulanix):

For this, you would use the pythagorean theorem:\[ a ^{2}+b ^{2}=c ^{2}\] Since the legs are 3 and 4, replace a and b with them. \[3^{2}+4^{2}=c ^{2}\] 9+16=c^2 25=c^2 Take the square root to get c=5

OpenStudy (anonymous):

so c is : \[c =\sqrt{3^2 + 4^2} =\sqrt{25} = 5\]

OpenStudy (anonymous):

^_^

OpenStudy (xkehaulanix):

Oh, sstarica has symbols. Use that one, it's prettier.

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

thanks that was very helpfull :)

OpenStudy (anonymous):

a pleasure ^_^